Question 26 (1 point) Calculate the market to book ratio of a firm with a current share price of $31.12, a book value of equity of $3,000,000,000, and 99,000,000 shares outstanding. Round your answer to four decimal places. Your Answer D View hint for Question 26 Question 27 (1 point) If a companys profit margin is 6% and its total asset turnover is 2.5, what is its return on assets (ROA)? (Ehter your answers as a decimal rounded to 4 decimal places, not a percentage. For example, enter 0.0843 instead Your Answer D View hint for Question 27
Answer #1

1. Book Value per share = 3,000,000,000/99,000,000 = 30.3030
Market to book Value = 31.12/30.3030 = 1.02696 or 1.0270

2. Profit margin = Net Profit/Sales = 6%
Total Asset turnover = Sales/Average Total Asset =2.5
ROA = Profit Margin* Total asset turnover = 6%*2.5 = 0.15

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well